Suppose that the demand and supply curves in the market for apples have the following functional forms: QD = 250 - 4p and QS = 10 + p. The equilibrium quantity and price would then be: ...Please explain.

A) Q = 58, p = 48

B) Q = 48, p = 58

C) Q = 68, p = 98

D) Q = 92, p = 48

E) Q = 68, p = 108
Answer #1

At the equilibrium the demand and supply are equal so we can write it as Qd= Qs

= 250 -4P = 10 + p

or 250 - 10 = P + 4p

= 240 = 5P

P = 240 / 5 = 48.

As the price is 48 we can put the number in any equation and get the value of the quantity.

250 - 4 (48) = 58. the answer is "A". price is 48 and the quantity is 58.